OpenStudy (anonymous):

write an equation for the line containing these points (2,2) and (3,-3)

OpenStudy (anonymous):

slope = -5/1 = -5 point slope form: y - 2 = -5(x-2) y-2 = -5x + 10 y = -5x + 12

OpenStudy (anonymous):

http://www.purplemath.com/modules/strtlneq2.htm